PRINTER VERSION 4 BAck NEXT Problem 2.064 Your answer is partially correct. Try again. In an acid-base titration, a base or acid is gradually added to the other until they have completely neutralized each other. Since acids and bases are usually colorless (as are the water and salt produced in the neutralization reaction), pH is measured to monitor the reaction. Suppose that the equivalence point is reached after approximately 100 mL of a NaOH solution have been added (enough to react with all the acetic acid present) but that replicates are equally likely to indicate from 95 to 104 mL to the nearest mL. Assume that volumes are measured to the nearest ml and describe the sample space. Round your answers to one decimal place (e.g. 98.7). a) What is the probability that equivalence is indicated at 100 mL? 0.1 b) What is the probability that equivalence is indicated at less than 100 mL? 0.5 c) What is the probability that equivalence is indicated between 99 and 100 mL (inclusive)? 0.1 Statistical Tables and Charts Click if you would like to Show Work for this question: Open Show Work

what is C's answer? zero?
Answer #1

Sample space will be 95,96,97,98,99,100,101,102,103,104

P(X =100) = 1/10= 0.1

b)

at least 100 will be 95,96,97,98,99,

P(X < 100) = 5/10= 0.5

c)

including 99 and 100 ml means 99 and 100

P(99<=X <=100) = 2/10=0.2